3 ms·
> Despite the fact that the number of students taking A-level maths has risen in recent years and that girls outperform boys at GCSE, the number of girls taking
by Dewie3 11y ago
> Despite the fact that the number of students taking A-level maths has risen in recent years and that girls outperform boys at GCSE, the number of girls taking A-level mathematics is proportionally much lower.
Grab the nearest popcorn and watch the gender wars unfold.
- escherplex 11y agoNot if this is considered a particularly difficult UK Maths GCSE exam question, which over 90 percent couldn't answer: Hannah has 6 orange sweets and some yellow sweets. Overall, she has n sweets. The probability of her taking 2 orange sweets is 1/3. Prove that: n^2-n-90=0. If a HS student I would imagine you would First: think of coins p(H1) = .5; p(H2) =.5; p(H1+H2) = .5 * .5 = .25 Second: OK, here (6/n) * (5/(n-1)) = 1/3 Third solve: 30/(n * (n-1)) = 1/3 90 = n^2 - n 0 = n^2 - n - 90 n=10 or -9; if -9 then Hannah only has one sweet and lifted 9 from somebody else just to demo her point. (from London Telegraph today)
- thaumasiotes 11y agomy instinct for approaching the problem: C(6,2) / C(n,2) = 1/3 (where C(n,k) is n choose k) 15 / [n(n-1)/2] = 1/3 30 / n(n-1) = 1/3 n(n-1) = 90 I probably wouldn't have done that in high school, though.